OpenStudy (anonymous):

a rectangle table is five times as long as it is wide if the area is 245ft2, find the length and the width of the table width is ___? length is ___?

OpenStudy (anonymous):

help someone a rectangle table is five times as long as it is wide if the area is 245ft2, find the length and the width of the table width is ___? length is ___?

OpenStudy (anonymous):

you have 2 equations and 2 unknowns length x width = 245 ft^2 length = 5*width so you sub the second equation into the first to get (5* width) * width = 245 and solve for width 5*width^2 = 245 width^2 = 245/5 width = sqrt(49) width = 7 ft now you sub that back into the second equation and get length = 5 * 7 length = 35 ft